### v1

11institutetext: Department of Mechanical Engineering, Texas A&\&&M University
11email: {kasiv, ssaripalli}@tamu.edu
22institutetext: DEVCOM Army Research laboratory, Adelphi, MD, USA
22email: {felix.a.sanchez12.ctr, timothy.j.overbye.ctr, jason.m.gregory1.civ}@army.mil
Trailblazer: Learning offroad costmaps for long range planning
Kasi Viswanath 11    Felix Sanchez 22    Timothy Overbye 22    Jason M. Gregory 22    Srikanth Saripalli 11
1 Introduction
Unmanned Ground Vehicles (UGVs) are crucial across missions such as search and rescue, environmental exploration, and surveillance in challenging off-road terrains. While recent advancements have enabled autonomous systems to achieve long-range navigation in structured on-road environments, off-road mobility remains constrained by dynamically changing terrain properties, unexplored pathways, and sensor uncertainty.
Though onboard sensors allow basic navigation, pre-loaded environmental data significantly boosts both safety and efficiency. This knowledge helps systems identify and avoid risky areas—such as unstable ground or steep slopes—while prioritizing safer routes. Without this understanding, robots rely solely on real-time sensing, increasing the likelihood of errors such as selecting unsafe paths. Integrating prior environmental information enables smarter, more reliable decision-making in unpredictable outdoor settings.
Overhead terrain data is a valuable source of prior environmental knowledge, especially when a vehicle encounters a site for the first time. These data sources include aerial or satellite imagery, digital elevation maps, and 3D LiDAR scans. Typically, this data is processed into 2D feature grids, which are then converted into traversal costs to guide autonomous navigation. While this approach has been effective in enhancing vehicle planning [1][2], it traditionally relies on hand-tuned cost functions, which are prone to bias.
Recent advancements have introduced imitation learning [3][4] and Vision Language Models [5] to optimize cost-maps through human demonstrations, reducing manual intervention. However, many of these methods still depend on pre-trained convolutional neural networks (CNNs) trained on manually crafted cost-maps, further refined using reinforcement learning algorithms.
Trailblazer offers a novel alternative by directly converting multi-modal sensor data into cost-maps. Instead of relying on pre-trained models or manual annotations, Trailblazer learns cost-map generation from expert driven demonstrations. This approach simplifies the process, eliminates manual tuning, and enhances the system’s adaptability and efficiency in real-world scenarios.
2 Methodology
Trailblazer methodology is divided into two key stages: (i) data collection and pre-processing, where satellite imagery and LiDAR data are gathered and prepared; (ii) Trailblazer framework and training, which involves the application of a sophisticated architecture for feature extraction and path planning.
2.1 Data collection and pre-processing
The Trailblazer architecture utilizes satellite/aerial imagery and airborne LiDAR point cloud data as primary inputs. This study sources data predominantly from the United States Geological Survey (USGS) and the Environmental Systems Research Institute (ESRI) via ArcGIS servers [6]. ESRI provides high-resolution satellite imagery with a spatial resolution of up to 30 cm per pixel, while the USGS LiDAR Explorer delivers point cloud data with a density exceeding 8 points per square meter. Imagery data contributes essential information about land cover, terrain, transportation networks, and water bodies, while LiDAR data offers detailed insights into terrain features such as slope profiles, height measurements, and intensity maps.
From the satellite image and aerial LiDAR data, we compute four maps as input for Trailblazer. The four maps consists of (i) Semantic segmentation, (ii) Height map, (iii) Slope profile (iv) Intensity map as shown in Figure 1.
Figure 1: Inputs to Trailblazer from Texas A&M RELLIS test site.
2.1.1 Segformer
SegFormer [7], a lightweight transformer-based architecture for semantic segmentation, is utilized to generate a 2D pixel-wise mask of the overhead image, classifying each pixel into distinct semantic classes. Segformer features a hierarchically structured transformer encoder coupled with a lightweight MLP decoder that effectively integrates both local and global attentions.
To train the SegFormer model, we use the FLAIR-one dataset [8], which provides semantic annotations for aerial imagery along with topographic and land cover information. This dataset consists of 77,412 high-resolution patches, each measuring 512×512 pixels with a spatial resolution of 0.2 meters, spanning 19 distinct semantic classes.
Category FLAIR Labels Class IoU (%)
Traversable Herbaceous region, Agricultural, Plowed land, Bare soil 71.88
Non-Traversable Pervious surface, Impervious surface, Ligneous, Mixed 52.96
Vegetation Coniferous, Deciduous, Brushwood, Vineyard, Clear cut 73.25
Obstacles Building, Greenhouse, Other 77.88
Water Water, Swimming pool 82.50
Mean IoU (mIoU) 71.69
Table 1: FLAIR Label Categories with Class IoU Metrics
We preprocess the FLAIR-one annotations by consolidating the 19 semantic classes into 5 superclasses, tailored to off-road terrains where fewer distinguishing features are typically required. This grouping is based on the semantic relevance of the classes within the environmental context. The resulting superclasses are outlined in Table 1.
2.2 Trailblazer Architecture
The Trailblazer architecture comprises two primary components: an encoder-decoder framework and a differential A* planner. The encoder-decoder architecture utilizes fully convolutional layers in conjunction with attention mechanisms to effectively extract features from the input data.
2.2.1 Encoder-Decoder:
Inspired by multiscale convolutional networks [9], Trailblazer integrates feature extraction and fusion modules. The feature extraction module comprises four convolutional layers and a pooling layer (Figure 2), enabling effective low-level feature sharing. To retain both translationally variant and invariant features, a spatial attention module [10] is incorporated at the skip connection. The module’s output is upsampled and concatenated with the skip connection, allowing independent treatment of feature channels. A final convolution on the concatenated output produces the cost map.
Figure 2: Trailblazer framework architecture. The whole network is differentiable.
2.2.2 Neural Astar Planner:
We integrate the Neural A* algorithm [11] into our neural network architecture to compute paths between start and goal points on a cost map. Neural A* is a data-driven search method that redefines the traditional A* algorithm to be differentiable, enabling seamless integration with a convolutional encoder to form an end-to-end trainable neural network planner. The differentiable A* module employs techniques like discretized activation inspired by Hubara et al. [12], allowing it to execute A* searches during the forward pass and back-propagate losses through all search steps to other trainable backbone modules. The loss function used is the mean L1 loss between the search history and the ground-truth path. This integration enhances the network’s ability to learn from expert demonstrations, producing paths that closely align with ground-truth trajectories in terms of accuracy and efficiency.
3 Experiments
3.1 Semantic Segmentation
We trained the Segformer model on the FLAIR-One dataset using a modified superclass annotation strategy to improve segmentation performance. By grouping classes into superclasses, we effectively reduced the complexity of the confusion matrix, leading to improved accuracy. The dataset was divided into 61,710 images for training and 15,700 images for validation. The model was trained for 400 epochs using a categorical cross-entropy loss function. This training process resulted in the model achieving a mIoU of 87.5% and an accuracy of 71.69% on the validation set. Examples of predictions on the validation set are illustrated in Figure 4.
3.2 Trailblazer
The Trailblazer model was initially trained using simulation data and later evaluated on real-world trajectories. For data collection, we employed the MAVS off-road simulator [13] to navigate diverse scenes and generate trajectory data. A total of 4,000 data instances were collected, comprising input maps and their corresponding trajectories. This dataset was split into training and validation subsets to train the Trailblazer model effectively. The training and validation loss curves, shown in Figure 4, demonstrate that the model achieved convergence, with a minimum validation loss of 0.0256. When tested on a separate dataset, the model achieved a loss of 0.0213, indicating its effectiveness in trajectory prediction.
Figure 3: Semantic ground truths and predictions from Segformer.
Figure 4: Training and validation loss curves of Trailblazer.
To enable sim-to-real transfer, we retrained the simulation-trained Trailblazer model using real-world trajectory data. The model was evaluated on various real-world overhead datasets obtained from the USGS, with an example illustrated in Figure 5.
3.2.1 Planned experiments:
OpenStreetMap (OSM) [14] provides GPS trajectories of vehicles driven across various locations, which we leverage to further train Trailblazer for enhanced generalization across diverse geographic regions. These trajectories, exemplified in Figure 2, serve as valuable data to improve the model’s adaptability. Concurrently, real-time navigation experiments are being conducted, where Trailblazer functions as a global planner in conjunction with a local planner. This integrated framework aims to enhance navigation efficiency and reliability, ensuring seamless movement between points in diverse environments.
Figure 5: Costmap, paths extracted from costmap and OpenStreetmaps tracks from a test site.
4 Experimental Insights
Traditional path planning relied on manual data collection via UAVs, limiting use in controlled/ restricted air spaces. Our work shows open-source data (USGS, ESRI, OpenStreetMap) enables robust urban/off-road navigation without costly manual efforts. Trailblazer achieves cross-terrain adaptability using these datasets, proving open-source data is sufficient for scalable global planning.
We experimented with replacing 3D LiDAR scans with Digital Elevation Maps (DEMs), which are widely available. While DEMs produced costmaps of comparable quality to 3D LiDAR, their effectiveness depended heavily on resolution. DEMs with resolutions below 1/3 arc-second resulted in subpar costmaps and slope distortions.
Trailblazer, initially a local planning algorithm for off-road terrains, now supports both local and global planning by dynamically updating the global costmap with real-time onboard data. For instance, if overhead data indicates a dry, traversable riverbed but local inspection reveals otherwise, the global costmap adapts, enabling a new optimal trajectory and ensuring robust planning in dynamic environments.